OpenStudy (anonymous):

Find the product of (u+4)(u^2-3u+3)

OpenStudy (whpalmer4):

Use the distributive property to multiply (this is why I hate that people teach FOIL!) \[(u+4)(u^2-3u+3) = u(u^2-3u+3) + 4(u^2-3u+3) = \]